OpenStudy (anonymous):

Solve the following systems using the substitution method. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.) 4x+3y=-4 y=3x-4

OpenStudy (anonymous):

First, you must get one variable to cancel out, so you must get either x or y to cancel out. 4x + 3y = -4 y = 3x-4 Due to the commutative property, this can be written as... 3y = -4x - 4 y = 3x - 4 Dividing the first equation by 3... y = -4/3x -4/3 y = 3x - 4 -------------------- Subtracting each from another y = -4/3x -4/3 -(y = 3x - 4) -------------- (-13/3x +8/3) = 0 *3 *3 -------------------- -13x + 8 = 0 -13x = -8 x = 8/13 And, pluggng this back into our original equations, y = - 28/13